In a certain species of fish green scales are dominant to red scales, and the presence of whiskers are dominant to the absence of whiskers. Both traits are autosomal and they are unlinked.

The frequencies of alleles for scales are; G=0.7 and g=0.3.

The frequencies of alleles for whiskers are ; W = 0.9 and w=0.1.

What is the expected frequencies for fish that have red scales and whiskers?

a. .089
b. .078
c. .065
d. .081
e. .073
Answer #1

Frequency of fish that have red scale = q2 = 0.3*0.3 = 0.09

Frequency of the fish that have whiskers are = p2 +2pq = (0.9*0.9) + 2* 0.9 * 0.1 = 0.81 + 0.81

Expected frequency of fish to have the red scale and whisker = frequency of fish to have red scale * frequency of fish to have whisker

= 0.09 * 0.99 = 0.0891 ~ 0.089

So the option first is correct